Accelerator Update for April 13, 2009 - April 15, 2009 |
Monday April 13 The day shift began with Operations monitoring the Tevatron (TeV) store 6974 (12.76E31) and the Recycler stash (197.1E10), with the Antiproton Source stacking (17.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. A Linac low energy RF station (LRF3) developed a LCW leak inside the modulator. Linac and Mechanical Support personnel repaired the station. LRF3 was operational by 7:45 PM. MTest resumed taking beam at 7:45 PM. Operations terminated store 6974 at 8:30 PM and followed with HEP shot setup. Operations established store 6975 by 10:00 PM with an average initial luminosity of 31.9E31. Operations disabled the MTest beam at 10:00 PM. Tuesday April 14 The midnight shift began with Operations monitoring store 6975 (20.6E31) and the stash (107.8E10), with Pbar stacking (11.8E10), and with NuMI and MiniBooNE taking beam. An AP-2 transfer line quadrupole power supply (D:Q731) faulted off at 12:40 AM. Pbar power supply techs returned D:Q731 to service by 1:45 AM. MTest began taking beam at 4:25 AM. Repairs to the Accumulator bend bus power supply, A:IB, began at 7:00 AM, once all antiprotons were transferred to the Recycler. The MTest T953 experiment completed their run shortly after the day shift began. The A:IB power cable repairs continued throughout the day shift. Pbar began stacking by 12:30 PM. Pbar transfers to the Recycler followed normally. Operations terminated store 6975 at 3:00 PM. CDF experimenters accessed their collision hall. A:IB developed an LCW leak towards the end of the day shift. CDF came out of access at 4:30 PM. A:IB repairs were completed by 4:40 PM and stacking resumed. HEP shot setup began at 6:15 PM and store 6978 was colliding by 7:30 PM with an average initial luminosity of 30.2E31. Wednesday April 15 The midnight shift began with Operations monitoring store 6978 (14.2E31) and the stash (141.6E10), with Pbar stacking (33.3E10), and with NuMI and MiniBooNE taking beam. At 12:58 AM, MI-60 suffered a vacuum burst. At 5:19 AM, Pbar stopped stacking. Operations discovered and fixed an energy mismatch between the Pbar Debuncher and the Accumulator. At 6:11 AM, the Pbar sweeper magnet tripped off. It reset. The Neutron Therapy Facility reported at 7:32 AM that they were preparing to treat their first of two patients for the day. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 04/15/2009 email Fermilab |
| Security, Privacy, Legal | |